### v1

# AffordDexGrasp: Open-set Language-guided Dexterous Grasp with Generalizable-Instructive Affordance

Abstract
Language-guided robot dexterous generation enables robots to grasp and manipulate objects based on human commands. However, previous data-driven methods are hard to understand intention and execute grasping with unseen categories in the open set. In this work, we explore a new task, Open-set Language-guided Dexterous Grasp, and find that the main challenge is the huge gap between high-level human language semantics and low-level robot actions. To solve this problem, we propose an Affordance Dexterous Grasp (AffordDexGrasp) framework, with the insight of bridging the gap with a new generalizable-instructive affordance representation. This affordance can generalize to unseen categories by leveraging the object’s local structure and category-agnostic semantic attributes, thereby effectively guiding dexterous grasp generation. Built upon the affordance, our framework introduces Affordacne Flow Matching (AFM) for affordance generation with language as input, and Grasp Flow Matching (GFM) for generating dexterous grasp with affordance as input. To evaluate our framework, we build an open-set table-top language-guided dexterous grasp dataset. Extensive experiments in the simulation and real worlds show that our framework surpasses all previous methods in open-set generalization.
{strip}[Uncaptioned image]
Figure 1: Open-set Language-guided Dexterous Grasp. Our framework bridges the gap between language and grasp actions through Generalizable-Instructive Affordance, which enables cross-category generalization via category-agnostic cues and graspable local structure. Remarkably, our framework demonstrates strong generalization without requiring extra real training data in real-world experiments.
1 Introduction
Achieving generalizable robot dexterous grasping is an important goal in the fields of robotics and computer vision, with exciting potential applications in human-robot interaction and robot manipulation.
Recent works explore the task of language-guided dexterous grasp generation[51, 20, 21, 59], aiming to enable dexterous hands to perform actions based on language instructions, going beyond previous works that focus on stable grasping [46, 54, 26, 38]. The typical language-guided approaches employ language as the condition of generative models to predict hand parameters, achieving impressive performance on objects within known categories [51, 4]. However, in the open real world, there are many categories that may not appear during training, and the cost of collecting data for dexterous hands is quite expensive. Therefore, the open-set generalization on unseen category samples is crucial for robot grasp. While previous works on parallel grippers explore open-set tasks [36, 23], there is limited research on open-set generalization for dexterous hands, as their significantly higher degrees of freedom present complex challenges.
In this work, we explore a novel and challenging task: open-set language-guided dexterous grasping, as shown in Figure 1, where models are evaluated on objects and language instructions from both seen and unseen categories. This task poses a great challenge in ensuring that grasps are intention consistent with corresponding language instructions for unseen categories. We find that the main challenge lies in the huge gap between high-level natural language and low-level robot action spaces, which makes it difficult to generalize the ability of understanding intentions and grasping from the training domain to unseen categories.
To solve the above challenge, we propose the Affordance Dexterous Grasp (AffordDexGrasp) framework, with the key insight of employing a new affordance as an intermediate representation to bridge the gap between high-level language and low-level grasp actions. Our affordance representation combines two characteristics: generalizable (generalizing to unseen categories based on language) and instructive (guiding dexterous grasp generation effectively). However, achieving these two characteristics is not trivial. For example, as shown in Figure 2, fine-grained contact information can effectively guide grasp generation or optimization [2, 9, 17, 55], but it is difficult to generalize to unseen categories. In contrast, coarse information, such as object parts, can be obtained from a pre-trained computer vision model [27, 43, 41], but it is too coarse to guide dexterous hand actions with higher degrees of freedom.
To achieve these characteristics, we propose the Generalizable-Instructive Affordance by defining a general graspable affordance that aligns with category-independent information, such as intention, object parts, and direction. As shown in Figure 2, the affordance represents the potential graspable regions of all grasps with the same semantics. In this way, the models do not need to learn complex dexterous contact patterns but instead focus on a general graspable area, which can be well aligned with category-agnostic semantic attributes and guide grasp generation effectively.
To generate affordance and employ it to guide grasp generation, our framework consists of two cascaded generative models. The Affordance Flow Matching generates affordance maps in a generalizable manner based on language. And the Grasp Flow Matching generates dexterous grasp poses under the effective guidance of affordance. Moreover, we introduce a pre-understanding stage and a post-optimization stage to further boost generalization. Specifically, we employ the Multimodal Large Language Model (MLLM) to pre-understand user intention to enhance generalization across diverse user commands. And we introduce an affordance-guided optimization to improve grasp quality while preserving consistency with user intentions.
For evaluating our framework, we build a open-set table-top language-guided dexterous grasp dataset, based on the language guided dexterous grasp dataset [51, 56]. We exclude specific categories from the training set to test the model’s open-set generalization. Moreover, we provide high-quality rendered images to facilitate the usage by MLLM, and we extend the dataset to scene-level data to better simulate real-world environments. The comprehensive experiments are conducted on both the simulation open-set dataset and real-world environments. The results show that our framework can generate dexterous grasp with consistent intention and high quality in open set.
2 Related work
2.1 Dexterous Grasp
Dexterous grasping is critical in robotics, which equips robots with human-like grasping capabilities. Some methods [29, 58, 44, 48] focus on grasp stability and quality, while recent studies explore task-oriented [50, 61] or language-guided grasping [21, 51] with specific semantic intention. For language-guided task, the data-driven methods achieve impressive performance in unseen object within seen categories by leveraging language conditioned generative model [4, 51]. However, we find that these models are hard to generalize for samples from unseen categories. And we find that the main challenge is the huge gap between language and grasp action. To solve this problems, we propose AffordDexGrasp framework with the insight that bridge the gap by generalizable-instructive affordance.
Figure 2: Different affordance representations. (a) While conatct map are too elaborate to generalize and object part are too coarse to guide grasping, our affrodacne achieve a balance. (b) While object part has a lower upper bound and contact shows significant degradation in generalization, only our affordance effectively achieve the balance (Top-1 indicates grasp intention consistency).
2.2 Open-set Robot Grasp
Exploring the performance of robotic models in open-set scenarios is crucial due to the diverse object categories in the real world and the high cost of data collection. For tasks that only focus on stable grasping, both parallel grippers [6, 52] and dexterous hands [47, 58] achieve impressive performance on open set. However, when it comes to considering task-oriented [45] or language-guided grasping [33], it becomes much more difficult. Some works in parallel grippers achieve this by using pre-trained visual grounding models [22, 41, 27], implicit feature semantic fields [18, 36, 32], or multimodal large models [13, 14]. Compared to parallel grippers, dexterous hands have a significantly higher number of degrees of freedom, making the methods of parallel grippers difficult to transfer to dexterous hand [51]. In this paper, we explore the open-set language-guided dexterous grasping task and propose a novel framework to address it.
2.3 Grasp Affordance
Affordance is first introduced in [8], referring to environmental action possibilities. There are currently two types of grasp affordance: fine-grained contact [42, 3] and coarse-grained object segmentation [27]. The fine-grained information, such as contact areas, normals, or key points, is commonly used in hand-object interaction[59, 62, 9, 2]. However, this fine-grained information is difficult to generalize well to unseen categories. As a result, inaccurate contact information would lead to unreasonable grasping. On the other hand, We find through experiments that the coarse object segmentation [27] is too coarse to guide dexterous grasp generation with relatively high degrees of freedom. To solve this problem, we propose Generalizable-Instructive Affordance, which can generalize through the local structure of objects using category-independent clues and effectively guide grasp generation.
Figure 3: The pipeline of Affordance Dexterous Grasp framework. The inference pipeline includes three stages: 1) intention pre-understanding assisted by MLLM; 2) affordance flow matching for generating affordance base on MLLM ouput; 3) Grasp Flow Matching and Optimization for outputing grasp poses based on the affordance and MLLM outputs. In the training time, AFM and GFM are independently trained one after another. Transformer and Perceiver are attention-based interaction module for velocity vector field prediction.
3 Affordance Dexterous Grasp Framework
Given the scene point clouds $\mathcal{O}$, RGB images $\mathcal{I}$ and language guidance $\mathcal{L}$ as input, our goal is to generate intention aligned and high-quality dexterous grasps poses $\mathcal{G}^{dex}$ = $(r,t,q)$, where $r$ denotes the rotation, $t$ denotes the translation, and $q$ represents the joint angles of the dexterous hand.
In this section, we first introduce a novel generalizable-instructive affordance representation (Section 3.1). Built on this affordance representation, we then propose the Affordance Dexterous Grasp framework (Section 3.2) including intention pre-understanding stage, Affordance and Grasp Flow Matching, and grasp post-optimization stage.
3.1 Generalizable-Instructive Affordance
We propose the generalizable-instructive affordance as an intermediate representation to bridge the gap between high-level language and low-level grasp actions. The affordances serves two primary objectives: 1) enable generalization by aligning object local structures with category-agnostic language semantics; and 2) provide instructive guidance for grasp generation through object affordance cues. However, there exists a trade-off in the generalization and instruction, and it is not trival to achieve both objectives.
Generalization vs. Instruction. While more fine-grained contact information can better guide grasp generation, it becomes harder to generalize to unseen categories. Fine-grained affordances typically compute the distance between object points and predefined elements, such as hand surface points or hand key points [28, 19]. However, we observe that these fine-gained representations are difficult to generalize to unseen category samples as shown in Figure 2. As the dexterous hand structure is complex and contact modes are varied, the contact maps are also elaborate and varies in space, which make it difficult to generalize.
On the other hand, the information agnostic to dexterous hands, like object part segmentation, can be obtained in a generalizable manner by vision models [27, 60] or semantic feature fields [36]. However, the object part is too coarse for dexterous hand, as it may struggle to define the clear grasping area. For example, grasping the body of a mug from above and from the side involves the same part, but with different intentions. Moreover, this coarse-grained learning may hinder the model from learning an accurate grasping method during training.
Generalizable-Instructive Affordance. To address this challenge, we propose a novel affordance representation that simultaneously enables both generalization and instruction. The key insight is to establish a correspondence between the general graspable affordance of objects and category-agnostic semantic attributes. Unlike contact maps that describe a particular hand contact area, our affordance represents all potential graspable regions that share the same semantic intention. This approach allows the affordance to generalize more effectively with category-agnostic semantic attributes from language, as it doesn’t have to learn the complex distribution of hand contact. Meanwhile, our affordance serves as priors to enhance grasp generation by providing valuable graspable area cues.
To obtain the ground truth of generalizable-instructive affordance, we first organize those grasps of one object with the same intention, contact parts and grasp direction into a semantic group. Then we calculate the distances between each scene points $\{o_{i}\}^{M}$ and hand surface points $\{h_{i}\}^{N}$, to obtain the contact map $\Omega_{contact}=\{d_{i}\}_{1}^{M}$ for each grasp in the group, where $d_{i}=\min\limits_{j=1\dots N}\left\lVert o_{i}-h_{j}\right\rVert_{2}$. Then we unite contact areas of $k$ grasps in each group by calculating the minimum value of contact areas: ${d^{min}_{i}}=\{\min\limits_{k}(d_{i}^{k})\}^{M}$. In order to make the affordance map more smooth, we apply Gaussian filtering to the union contact map: $a_{i}=\sum_{j\in\mathcal{N}(i)}\hat{w}_{ij}d_{j}^{min}$ and the weights are obtained by:
$\hat{w}_{ij}=\frac{\exp\left(-\|o_{i}-o_{j}\|^{2}/(2\sigma^{2})\right)}{\sum_{k\in\mathcal{N}(i)}\exp\left(-\|o_{i}-o_{k}\|^{2}/(2\sigma^{2})\right)},$ (1)
where $\mathcal{N}(i)$ is the neighborhood of $o_{i}$, $\sigma$ is set as the average nearest neighbor distance. Finally, we obtain the ground truth of generalizable-instructive affordance $\Omega=\{a_{i}\}^{M}$.
3.2 AffordDexGrasp Framework
3.2.1 Intention Pre-Understanding
The variability of user expression in the open world presents a challenge for models in generalizing and understanding user intentions. To address this, we employ the Multimodal Large Language Model (MLLM) to pre-understand user intentions, by imputing the user language and the rendered RGB image into GPT-4o [15] for intention understanding and key cues extraction. We empirically find that the following cues are both helpful for grasping and easily reasoned by the MLLM: the object category, user intention (e.g., use it), the contacting part of the object (e.g., the handle of a mug), and the grasp direction. The key cues can be extract from language when the commands are clear, otherwise can be reasoned from the image.
To improve the understanding ability of MLLM, we employ chain of thought (CoT) and visual prompt, and the prompts can be found in supplementary material A.4. Additionally, the grasp direction is simplified as a discretized direction in six predefined coordinates (front, back, left, right, up, down). We first ask MLLM to obtain the direction in image coordinates, and transform it to world coordinates by the camera pose. Then the direction vector is derived from the index of the nearest coordinate axis to form discretized direction. Finally, MLLM organizes this information into a concise sentence (e.g.,“use the mug from the left by contacting the handle”). This compact and information-dense sentence structure allows subsequent models to better understand users’ intentions.
3.2.2 Affordance Flow Matching
We introduce Affordance Flow Matching (AFM) to learn the affordance distribution efficiently, achieving intention aligned and generalizable affordance generation. AFM is built on the conditional flow matching model [30, 1, 25] aiming to learn the velocity vector field from the noised affordance map $\Omega_{0}$ to the target affordance map $\Omega_{1}$, and the objective loss of AFM could be:
$\mathcal{L}_{AFM}=\mathbb{E}_{t,\Omega_{0}\sim p_{0},\Omega_{1}\sim p_{1}}\left\|v_{\eta}(\Omega_{t},t|c)-(\Omega_{1}-\Omega_{0})\right\|^{2},$ (2)
where $p_{0}$ is Gaussian distribution with zero mean and unit variance at time $t=0$, and $p_{1}$ is the target distribution as time $t=1$. $v_{\eta}(\Omega_{t},t|c)$ is parameterized as a neural network with weights $\eta$ and $c$ is the input condition features.
Specifically, the input scene object point cloud $\mathcal{O}$ is encoded by Pointnet++ [34], language is encoded by a pre-trained CLIP model [35], and the direction vector and time embedding are encoded by the Multilayer Perceptrons (MLP). Due to the permutation invariance of point level affordance, the noisy affordance are concatenated with object features to aware the position and sematic information. Then all features are fed into a Perceiver IO Transformer, following [16, 49]. The noisy affordance features are served as input and output array, and the features of language and direction are served as latent array. The outputs of the decoder go through an MLP to obtain the prediction of the flow $v_{\eta}(\Omega_{t},t|c)\in R^{M}$. Finally, in the inference time, we adopt mulitple steps $1/\Delta t$ following the equation:
$\hat{\Omega}_{t+\Delta t}=\Omega_{t}+\Delta t\,v_{\eta}(\Omega_{t},t|c_{1}),\quad\ t\in[0,1).$ (3)
3.2.3 Grasp Flow Matching
We introduce Grasp Flow Matching (GFM) to learn the grasp distribution, conditioned on the output of MLLM and AFM. GFM is also built upon the flow matching model, and the objective is to generate a parameterized dexterous hand pose $\mathcal{G}^{dex}$ = $(r,t,q)$, which represents the global rotation, translation, and joint angles. In the training time, grasp generation requires calculating explicit losses in 3D space, which involves obtaining the target dexterous poses, which go through forward kinematics to obtain hand configuration [53]. To obtain the target pose, we estimate it in one step during the training stage:
$\hat{\mathcal{G}^{dex}_{1}}=\mathcal{G}^{dex}_{t}+(1-t)*v_{\theta}(\mathcal{G}^{dex}_{t},t|c_{2}),\quad\ t\in[0,1),$ (4)
where the condition features $c_{2}$ are the concatenation of object affordance, language and direction features. And the object affordance features are obtained by feeding affordance with point clouds into PointNet++ [34].
The losses of grasp generation include 1) the grasp pose regression L2 loss; 2) the hand chamfer loss to minimize the discrepancies between the actual hand shapes; 3) the hand fingertip key point loss to minimize the distance between fingertip position. The loss function can be formulated as:
$\mathcal{L}_{GFM}=\lambda_{pose}\mathcal{L}_{pose}+\lambda_{chamfer}\mathcal{L}_{chamfer}+\lambda_{tip}\mathcal{L}_{tip},$ (5)
where $\lambda_{pose},\lambda_{chamfer}$ and $\lambda_{tip}$ are loss weights.
The object penetration loss, which is used to penalize the penetration of hands into objects, is excluded from grasp generation training, as it leads to significant challenges in the training process [51]. Employing the penetration loss can reduce object penetration, but negatively impacts both intention alignment and generation diversity. Therefore, to achieve semantic-aligned grasp generation, we exclude object penetration loss and introduce affordance-guided grasp optimization to reduce object penetration.
Intention Quality Diversity
$FID\downarrow$ $CD\downarrow$ $R-Precision\uparrow$ $Suc.\uparrow$ $Q1\uparrow$ $Pen.\downarrow$ $\delta_{t}\uparrow$ $\delta_{r}\uparrow$ $\delta_{q}\uparrow$
$Top1$ $Top2$ $Top3$
Open Set A
Contact2Gen[28] 0.428 9.38 0.164 0.232 0.289 11.6% 1.71e-4 1.09 2.86 18.4 39.2
Contact2Grasp[19] 0.426 10.2 0.257 0.337 0.398 16.5% 0.0172 0.668 5.73 4.84 48.3
GraspCVAE[40] 0.378 5.55 0.309 0.402 0.469 21.9% 0.0150 0.709 4.36 4.60 21.2
SceneDiffuser[12] 0.303 5.02 0.397 0.473 0.523 29.1% 0.0151 0.487 7.37 9.44 68.9
DexGYS[51] 0.297 4.63 0.317 0.401 0.463 44.2% 0.0512 0.362 6.30 6.79 54.7
Ours 0.231 3.81 0.480 0.588 0.666 45.1% 0.0531 0.293 7.48 6.98 61.5
Open Set B
Contact2Gen[28] 0.492 8.95 0.196 0.257 0.299 6.04% 1.30e-4 1.11 3.26 18.7 45.5
Contact2Grasp[19] 0.369 11.9 0.248 0.318 0.367 16.2% 0.00801 0.798 5.22 4.63 40.7
GraspCVAE[40] 0.365 5.35 0.297 0.357 0.403 24.2% 0.00421 0.738 3.96 5.47 25.6
SceneDiffuser[12] 0.271 6.50 0.302 0.355 0.393 30.7% 0.00851 0.683 5.41 6.13 72.9
DexGYS[51] 0.358 3.40 0.294 0.358 0.403 35.2% 0.0220 0.691 5.40 6.11 59.0
Ours 0.162 2.95 0.532 0.609 0.655 38.9% 0.0352 0.361 6.86 6.84 56.3
Table 1: Results on open-set datasets compared with the SOTA methods.
3.2.4 Affordance-guided Grasp Optimization
We introduce a non-parametric test-time adaptation (TTA), affordance-guided grasp optimization, to improve the grasp quality while maintaining intention consistency. Compared to learning-based TTA methods [17, 51], our non-parametric optimization does not suffer from performance degradation caused by out-of-domain open-set samples.
We design optimization objectives based on that high-quality grasps should exhibit high grasp stability and intention alignment in grasping posture and contact areas. Specifically, our optimization objectives includes: 1) affordance contact loss for constraining hand contact candidate points in contact with the affordance area:
$\mathcal{L}_{\text{aff-dist}}=\sum_{i}\mathbb{I}\left((d(p^{c}_{i})<\tau_{1})\lor(d(p^{r}_{i})<\tau_{1})\right)\cdot d(p^{r}_{i}),$ (6)
where $d(p)=\min_{o\in\{o_{j}\in\mathcal{O}\mid a_{j}>\tau_{2}\}}\|p-o_{j}\|_{2}$, and $o_{j}$ and $a_{j}$ represent the position and affordance values of object points, and $p^{c}_{i}$ and $p^{r}_{i}$ represent the hand contact candidates of the initial coarse hand and current refined hand. 2) affordance fingertip loss to keep the fingertip positions that are in contact with the affordance area unchanged.
3) object penetration loss for punish hand-object penetration, 4) self-penetration loss to punish the penetration of hand it self. 5) joint limitation loss to punish out-of-limit joint angles. The optimization objectives can be formulated as:
$\min_{\mathcal{G}^{dex}}(\lambda_{1}\mathcal{L}_{\text{aff-dist}}+\lambda_{2}\mathcal{L}_{\text{aff-finger}}+\lambda_{3}\mathcal{L}_{pen}+\lambda_{4}\mathcal{L}_{spen}+\lambda_{5}\mathcal{L}_{joint}).$ (7)
4 Open-set Dexterous Grasp Dataset
To support the evaluation of our framework, we build an open-set dexterous grasp dataset in table-top environment based on the language-guided dexterous grasp dataset [51, 56]. Our open-set dataset consists of 33 categories, 1536 objects, 1909 scenes, and 43,504 dexterous grasps for Shadow Hand [37] and Leap Hand [39].
- Open Set Data Split. We divide the categories in our dataset into two groups, A and B, excluding one from the training set each time to form Open Set A and B. Open Set A consists of 9,688 samples of 808 objects from 10 unseen categories, and 159 unseen objects from seen categories. Open Set B contains 29,744 samples of 568 unseen objects from 10 unseen categories, and 202 unseen objects from seen categories.
- Scene Construction. To make our dataset more practical, we generate scene data by placing objects in a tabletop environment using Blenderproc [5]. To prevent collisions between the hand and the table, objects are elevated using a shelf, and collision detection is performed to filter out invalid grasps. The physics engine is activated to ensure physically plausible object placements and grasping poses, enhancing dataset quality.
- Scene Point-Cloud and Image Rendering. The scene was captured using five RGB-D cameras: four positioned at elevated lateral angles and one directly above the object. The partial point clouds are merged into a global point cloud. To render realistic RGB images for MLLM processing, we use the texture generation model Paint3D [57] to apply realistic textures to all objects in our dataset.
Open Set A Open Set B
$FID\downarrow$ $CD\downarrow$ $Top1\uparrow$ $Q1\uparrow$ $Pen.\downarrow$ $FID\downarrow$ $CD\downarrow$ $Top1\uparrow$ $Q1\uparrow$ $Pen.\downarrow$
Necessity of Generalizable-Instructive Affordance
w/o affordance 0.338 5.09 0.369 0.0180 0.569 0.322 4.44 0.308 0.0194 0.452
w object part (pred) 0.313 5.06 0.320 0.0138 0.594 0.316 3.79 0.344 8.12e-3 0.721
w contact map (pred) 0.320 4.89 0.380 0.0131 0.605 0.286 3.25 0.408 9.25e-3 0.636
w our affordance (pred) 0.242 3.79 0.480 0.0240 0.501 0.176 2.76 0.538 0.0150 0.612
w object part (GT) 0.306 5.05 0.348 0.0198 0.523 0.271 3.79 0.353 6.68e-3 0.526
w contact map (GT) 0.113 2.01 0.598 0.0285 0.494 0.138 2.68 0.594 0.0143 0.643
w our affordance (GT) 0.169 3.52 0.509 0.0181 0.545 0.164 2.61 0.550 0.0120 0.584
Effectiveness of Pre-Understanding Stage
w/o key cues extraction 0.258 3.99 0.463 0.0222 0.510 0.186 3.42 0.515 8.33e-3 0.702
w/o direction 0.254 4.07 0.432 0.0162 0.531 0.185 2.85 0.529 8.80e-3 0.679
w MLLM 0.242 3.79 0.480 0.0240 0.501 0.176 2.76 0.538 0.0150 0.612
Effectiveness of Affordance-based Optimization
w/o our optimization 0.242 3.79 0.480 0.0240 0.501 0.176 2.76 0.538 0.0150 0.612
w penetration loss 0.336 7.07 0.318 0.0505 0.299 0.241 5.76 0.445 0.0147 0.719
w ConatctNet [17] 0.349 12.3 0.279 0.0434 0.124 0.355 11.1 0.240 0.0206 0.129
w RefineNet [51] 0.249 3.96 0.436 0.0487 0.399 0.203 2.92 0.493 0.0149 0.657
w our optimization 0.231 3.81 0.477 0.0531 0.293 0.162 2.95 0.532 0.0350 0.361
Table 2: Ablation study for our framework. The results of first two experiment groups are obtained from model outputs without optimization for an intuitive and reasonable comparison. The results in each group should be compared with light yellow line (our default setting).
5 Experiment
5.1 Evaluation Metrics
We employ three types of metric to evaluate the ability of intention consistency, grasp quality and grasp diversity.
- Evaluating Intention Consistency. We employ several metrics, including FID, R-Precision and Chamfer Distance. 1) FID (Frechet Inception Distance) [11] measures the distribution similarity between generated grasps and ground truth. To extract grasp and instruction features, we train an object-grasp point cloud encoder and a language encoder by contrastive learning [10]. 2) R-Precision evaluates the semantic alignment between language instructions and generated grasps [10]. Specifically, for each generated grasp, we construct a pool of 32 samples, including its ground-truth instruction and randomly selected samples. We then compute and rank the cosine distances between the point cloud features and the language features of all samples in the pool. The average accuracy is computed at the top-1, top-2, and top-3 positions, that a retrieval is considered successful if the ground truth entry appears among the top-k candidates. 3) Chamfer Distance (CD) quantifies the discrepancy between the predicted hand point cloud and the closest target with same intention [51].
- Evaluating Grasp Quality. Following [46], we use success rate, denoted as $Suc.$ , in Issac Gym [24] and Q1 [7] to assess grasp stability. Maximal penetration depth (cm), denoted as $Pen.$ is used to calculate the maximal penetration depth from the object point cloud to hand meshes.
- Evaluating Grasp Diversity. We compute the Standard deviation of translation $\delta_{t}$, rotation $\delta_{r}$ and joint angle $\delta_{q}$ of samples within the same scene observation.
5.2 Implementation Details
For training our framework, the number of epochs is set to 50 for AFM and 200 for GFM. For GFM, the loss weight of L2 loss is set to 1. For GFM, we set $\lambda_{\text{pose}}=10$, $\lambda_{\text{chamfer}}=1$, and $\lambda_{\text{tip}}=2$. The batch size is set to 16 for AFM and 64 for GFM. The initial learning rate is $2.0\times 10^{-4}$, decaying to $2.0\times 10^{-5}$ using a cosine learning rate scheduler [31]. The Adam optimizer is used with a weight decay rate of $5.0\times 10^{-6}$. In the inference, the time step is set to 10 for AFM and 20 for GFM. For the optimization, $\lambda_{1}=100$, $\lambda_{2}=10$, $\lambda_{3}=100$, $\lambda_{4}=10$, $\lambda_{5}=100$. The number of optimization iterations is set to 200. All experiments are implemented using PyTorch on a single RTX 4090 GPU.
5.3 Comparison with SOTA Methods
We compare our methods with the SOTA methods, as shown in Table 1. The generic grasp methods are reproduced by concatenating the point cloud features and features of same language guidance with ours. The same encoders are employed for fair comparison, and the penetration loss is excluded to avoid learning difficulties according to [51]. The results show that our framework significantly outperforms all previous methods in terms of intention consistency and grasp quality. Our framework also achieves a reasonably high level of diversity, as excessive diversity may lead to unnatural postures. Previous language-guided methods fail to generalize well to unseen categories due to the huge gap between language and grasping actions. Similarly, contact-based methods don’t perform well, as the contact maps are difficult to generalize. Additionally, our method outperforms other methods in the close set, as shown in Table 3. Overall, the results indicate that our framework achieves strong performance in intention-aligned, high-quality, and diverse grasps.
Figure 4: The visualization of generated affordance and dexterous grasp. The left top shows the zero-shot samples and the left bottom shows the one-shot samples in real world. The right top and right bottom show the zero-shot samples in simulation open set A and B.
$FID\downarrow$ $CD\downarrow$ $Top1\uparrow$ $Q1\uparrow$ $Pen.\downarrow$
GraspCAVE 0.208 2.40 0.395 0.0100 0.771
DexGYS 0.0804 1.74 0.590 0.0551 0.397
Ours 0.0286 1.13 0.779 0.0562 0.193
Table 3: Experiment of close set with SOTA methods.
5.4 Necessity of Our Affordance
The results shown in Table 2 validate the key insight of our framework: using the generalizable-instructive affordance to bridge the gap between high-level language and low-level grasp actions. The first row presents the results without affordance, which fails to generalize to unseen categories. The subsequent two rows show the results under the condition of object parts and contact maps. Prediction refers to maps generated by models, while GT refers to using the corresponding ground truth. As discussed in Section 3.1, the results demonstrate that object parts are too coarse to guide the grasp effectively, even when using the ground truth. On the other hand, the contact map is too fine-grained to generalize to unseen categories. Only our affordance achieves a balance between generalization and instruction.
Seen Cate Similar Cate Novel Cate
$Top1\uparrow$ $CD\downarrow$ $Top1\uparrow$ $CD\downarrow$ $Top1\uparrow$ $CD\downarrow$
Zero-Shot
w/o afford 0.827 2.61 0.295 5.71 0.219 7.82
Ours 0.880 1.06 0.432 3.65 0.246 6.42
One-Shot
w/o afford 0.767 3.17 0.429 3.67 0.342 6.64
Ours 0.870 0.99 0.656 1.73 0.586 3.03
Table 4: Zero-shot and one-shot generalization of our framework.
Furthermore, our affordance representation and framework demonstrate good generalization in the one-shot setting, as shown in Table 4. For the one-shot experiments, we introduce several novel categories to the test set that differ significantly from the training set and add one object from each unseen category to the training set. The results show that our framework achieves a significant performance improvement, not only on samples from different categories with similar structures but also for novel categories.
5.5 Effectiveness of Each Component
Further ablation studies are conducted to evaluate each component of our framework. For the pre-understanding stage assisted by MLLM, w/o key cues extraction refers to directly inputting the user’s commands into the model, while w/o direction means that the direction information is not utilized. Both designs enhance performance, and we believe that MLLM would be more beneficial in real-world applications due to its powerful generalization. For affordance-based optimization, our optimization improves grasp quality and maintains intention consistency, outperforming the learning-based methods ContactNet [17] and RefineNet [51], as well as the model which is trained with penetration loss.
5.6 Real World Experiments
The real-word experiments are conducted to verify the simulation-to-reality ability of our framework. We employ a Leap Hand, a Kinova Gen3 6DOF arms and an original wrist RGB-D camera of Kinova arm. And we collect several common objects in daily life, including unseen objects and unseen categories. In experiment, we synthesize the scene point cloud by taking several partial depth maps around the object. Then the scene point cloud, a RGB image and the user language command are fed into our framework to obtain the dexterous grasp pose. During execution, we first move the the arm to a pre-grasp position, then synchronously move the joints of the robotic arm and the dexterous hand to reach the target pose. In evaluation, a grasp is considered as success if the hand can lift it up and the action is consisted with the given command. The results shown in Table 5 shows that our framework achieve higher success rate than previous method.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
DexGYS 2 7 3 6 4 7 0 10
Ours 10 5 9 10 6 10 6 10
Table 5: Real world experiments: Successes in 10 attempts.
6 Conclusions
We believe that achieving generalizable dexterous grasps in open set is crucial within the deep learning and robotics communities. In this paper, we explore a novel task of open-set language-guided dexterous grasp. This task is challenging due to the huge gap between language and grasping actions, which hinders the model’s generalization ability. We propose the AffordDexGrasp framework with the insight that using a new affordance representation, generalizable-instructive affordance, to solve this challenge. This affordance enables generalization to unseen categories by utilizing the object’s local structure and category-agnostic semantic attributes, thus facilitating effective dexterous grasp generation. Based on this affordance, our framework introduces two flow matching based models for affordance generation and affordance-guided grasp generation. Moreover, we introduce a pre-understand stage and a pose-optimization stage to further boost generalization. We conduct extensive open-set experiments in both simulation and the real world, and the results show that our framework outperforms all SOTA methods.